# Faster Reinforcement Learning by Freezing Slow States

### v2

@@ 13       Sample basepoints $\mathcal{D}=\{(x^{(i)},y^{(i)})\}_{i=1}^{N}\sim\rho$
The agent wishes to maximize infinite horizon discounted reward, with $\gamma=0.995$. We omit a detailed formulation of the MDP for simplicity.
Figure 8: Mean return (with 95% CI) versus computational cost for the spatial gridworld problem across five independent replications.
The results are shown in Figure 8. In contrast to the inventory control problem, the performance differences between FSVI algorithms with different values of $T$ are far more pronounced in this domain: the smallest $T$ value of $T=3$ improves gradually and consistently, but is never quite able to achieve good performance. This effect is likely due to the nature of the tasks and the need to coordinate navigation across multiple steps. When $T$ is too small, the frozen lower-level policy is unable to fully capture meaningful navigation plans, leading to suboptimal behavior at the upper level. Interestingly, we observe that an intermediate value such as $T=6$ provides a strong compromise: it allows the lower-level policy to encode useful subtask behavior while retaining enough flexibility for the upper level to guide high-level decisions. We also notice that Base VI and QI slowly improve and eventually reach optimality, but requires nearly 3x the computation to reach 75% optimality that FSVI with $T=6$ is able to achieve very quickly.
@@
We consider a dynamic pricing problem where a seller interacts with a market of consumers who exhibit reference effects—that is, their purchase behavior depends not only on the current price but also on a history of past prices. Reference prices represent consumers’ internal price expectations for the product, based on historical prices. When the firm sets prices below (above) the consumer’s reference price, there is a positive (negative) impact on demand. This is motivated by real-world e-commerce and retail environments where consumers are sensitive to perceived fairness or price changes over time. The model we consider is a slight extension of the one in Chen et al. (2017).
@@
Figure 9: Mean return (with 95% CI) versus computational cost for the dynamic pricing problem using linear architectures across five independent replications.
Figure 10: Mean return (with 95% CI) versus computational cost for the dynamic pricing problem using linear network architectures across five independent replications.
As shown in Table 2, this is the largest problem out of the three. We therefore use this domain to test the “fitted” variants of our algorithms. Figures 9 and 10 compare the performance of frozen-state and baseline methods under two different function approximation regimes: linear (Figure 9) and neural network (Figure 10). We point out several takeaways:
@@
    Frozen-state methods improve upon baselines in both settings, but the gains are larger with neural networks. This is largely due to the relatively poor performance of Deep-FVI and Deep-FQI, which appear to suffer from higher variance and instability during training. We hypothesize that this is due to the difficulty of fitting high-capacity models under long-horizon dynamics, where bootstrapping can amplify noise. In contrast, the structured nature of Deep-FSFVI, specifically the reduced effective discount factor and frozen slow states, likely contributes to improved stability and more reliable learning.
@@
    Again, we see that that the $T=3$ version of FSVI performed poorly under both architectures, but $T=6$ and $T=12$ were both adequate.
@@ Conclusion
Appendix A Experimental Details
Architectures for FVI, FQI, and FSFVI
For the linear variants of these algorithms (Lin-FVI, Lin-FQI, and Lin-FSFVI), we use a linear regression architecture, mapping directly from the state vector to the target value (i.e., we did not transform the states into features). For the neural network variants of these algorithms (Deep-FVI, Deep-FQI, and Deep-FSFVI), we use a three-layer feedforward neural network consisting of two hidden layers with ReLU activations and a final linear output layer that maps to a scalar.
Inventory Control with Fixed Order Costs
In addition to the parameter setting given in the main paper, for each value function backup, we use 50 samples from the simulator for the upper level of FSVI, Base VI, QI, and Slow-agnostic VI. For the lower level of FSVI, we are in a deterministic setting and simply use 1 sample. In Figure 6, we plot roughly three data points per sweep of the state space (i.e., per single value iteration). To compute the optimal value, we ran Base VI for 50 iterations with 100 samples per backup.
@@ Appendix A Experimental Details
Dynamic Pricing with Reference Effects
Appendix B Value Function Approximation with a Linear Architecture
@@
It is well-known that naive specifications of approximate value iteration applied to linear architectures can produce divergent behavior (Bertsekas and Tsitsiklis, 1996b). To circumvent this potential issue, our algorithmic approach depends on a set of pre-selected states $\tilde{\mathcal{S}}=\{s_{1},s_{2},\ldots,s_{M}\}$, an idea popularized in Tsitsiklis and Van Roy (1996), who showed that if certain assumptions on these states and the feature vectors are satisfied, then divergence is avoided. A similar algorithm is also described more recently in Zanette et al. (2019).
@@
Appendix C Proofs from Sections 3 and 4
@@
Lemma C.1 is a consequence of Assumption 1. ∎
@@
Lemma C.3 (Discrepancy between $H$ and $\tilde{H}$).
@@
Proof of Proposition 6.1
Noting that $U^{*}$ has Lipschitz constant $L_{U}$ by Assumption 2, we can apply Lemma C.3 to Term A to obtain
Moving on to Term B, since the reward function $r\geq 0$, it follows that $U^{*}(s)\geq 0$ for all $s$. Also, the monotonicity of $\tilde{H}$ implies that $(\tilde{H}^{T-1}U^{*})\geq(\tilde{H}^{T-1}0)$. Therefore, applying Lemma C.2,
$=\gamma^{T-1}\textstyle\max_{y\in\mathcal{Y}_{s_{1}}^{T-1}}U^{*}(x_{1},y),$
$\leq\gamma^{T}L_{U}d_{\mathcal{Y}}(\alpha+2)(T-1),$ (30)
where (29) follows by Assumption 2 and (30) comes from Lemma C.1 (we use that $x_{T}(\omega)$ is $T-1$ transitions from $x_{1}(\omega)$ and both $\tilde{y}_{s_{1}}(\omega)$ and $y_{T}(\omega)$ are both $T-1$ transitions from $y_{1}(\omega)$). Finally, we have
[…]
which completes the proof.
Appendix D Proofs for Section 6.2
@@
Proof of Lemma 6.1
Let $\epsilon_{r}(\bm{\pi}^{*},J_{1})=\max_{s,a}\,|\mathbb{E}[R(s,a,\bm{\pi}^{*})]-\mathbb{E}[\tilde{R}(s,a,J_{1})]|$, so that we have $|r_{1}(s,a)-r_{2}(s,a)|\leq\epsilon_{r}(\bm{\pi}^{*},J_{1})$. Noting that the first steps of $f_{1}$ and $f_{2}$ are the same (action $a$ in state $s$ with $w_{1}$ revealed), applying parts (ii) and (iii) of Lemma C.1, the maximum discrepancy between $f_{1}$ and $f_{2}$ is:
Applying Lemma D.1, we see that
We also need to account for the fact that $\mu$ is greedy with respect to $V$, an approximation of the optimal value of $\mathcal{M}_{2}$. More precisely, $\mu$ is greedy with respect to $r_{2}(s,a)=\mathbb{E}\bigl{[}\tilde{R}(s_{0},a,J_{1})\bigr{]}$, $f_{2}(s,a,\bm{w})=s_{T}(a,\bm{\pi})$, and value function $V$. We can thus apply Lemma D.2 with $\epsilon_{r}=\epsilon_{r}(\bm{\pi}^{*},J_{1})$, $d=d(\alpha,d_{\mathcal{Y}},T)$, $L_{1}=L_{U}$, $\epsilon_{U}=\bigl{\|}\bar{U}^{*}-V^{*}(J_{1},\bm{\pi})\bigr{\|}_{\infty}$, and $\tilde{\epsilon}_{U}=\bigl{\|}V-V^{*}(J_{1},\tilde{\bm{\pi}})\bigr{\|}_{\infty}$. Collecting terms completes the proof.
Appendix E Proofs for Section 5
Technical Lemmas
@@ Appendix E Proofs for Section 5
Lemma E.2 (Proposition 6.1 of Bertsekas and Tsitsiklis (1996b)).
@@
Since $\nu_{k}$ is greedy with respect to $U_{k}$, we can apply Lemmas E.1 and E.2 to obtain
@@ Appendix F Bounds on $L_{U}$ in Terms of $L_{r}$ and $L_{f}$
Appendix G Proofs for Appendix B

### v3

@@ 11       Sample a dataset $\mathcal{D}=\{(x^{(i)},y^{(i)})\}_{i=1}^{N}\sim\rho$
The agent wishes to maximize infinite horizon discounted reward, with $\gamma=0.995$. Again, we use $M=M_{u}=50$ for all methods, and $M_{l}=1$ for E-FSVI. We omit a detailed formulation of the MDP for succinctness, but give the full reward specification in Appendix B.2.
Figure 8: Mean return (with 95% CI) versus computational cost for the spatial gridworld problem across five independent replications. The E-FSVI, we plot the performance of the policy after each lower-level value iteration and 2 times per upper-level value iteration. For the other methods, we plot 2 times per value iteration. To compute the optimal value, we ran Base E-VI for 50 iterations with $M=100$ samples per backup.
The results are shown in Figure 8. In contrast to the inventory control problem, the performance differences between E-FSVI algorithms with different values of $T$ are far more pronounced in this domain: the smallest $T$ value of $T=3$ improves gradually and consistently, but is never quite able to achieve good performance. This effect is likely due to the nature of the tasks and the need to coordinate navigation across multiple steps. When $T$ is too small, the frozen lower-level policy is unable to fully capture meaningful navigation plans, leading to suboptimal behavior at the upper level. Interestingly, we observe that an intermediate value such as $T=6$ provides a strong compromise: it allows the lower-level policy to encode useful subtask behavior while retaining enough flexibility for the upper level to guide high-level decisions. We also notice that Base VI and QI slowly improve and eventually reach optimality, but requires nearly 3x the computation to reach 75% optimality that E-FSVI with $T=6$ is able to achieve very quickly.
@@
We consider a dynamic pricing problem where a seller interacts with a market of consumers who exhibit reference effects: that is, their purchase behavior depends not only on the current price but also on a history of past prices. Reference prices represent consumers’ internal price expectations for the product, based on historical prices. When the firm sets prices below (above) the consumer’s reference price, there is a positive (negative) impact on demand. This is motivated by real-world e-commerce and retail environments where consumers are sensitive to perceived fairness or price changes over time. The model we consider is a slight extension of the one in Chen et al. (2017).
@@
Figure 9: Mean return (with 95% CI) versus computational cost for the dynamic pricing problem using linear architectures across five independent replications. The E-FSVI, we plot the performance of the policy after each lower-level value iteration and once per upper-level value iteration. For the other methods, we plot once per value iteration.
Figure 10: Mean return (with 95% CI) versus computational cost for the dynamic pricing problem using linear network architectures across five independent replications. The E-FSVI, we plot the performance of the policy after each lower-level value iteration and once per upper-level value iteration. For the other methods, we plot once per value iteration.
As shown in Table 2, this is the largest problem out of the three. We therefore use this domain to test the “fitted” variants of our algorithms. We set $N$, the number of state samples, to be 10% of the size of the state space. As before, we use $M=M_{u}=50$ for all methods, and $M_{l}=1$ for E-FSVI.
Figures 9 and 10 compare the performance of frozen-state and baseline methods under two different architectures: linear function approximation (Figure 9) and neural network function approximation (Figure 10). We point out several takeaways:
@@
    Frozen-state methods improve upon baselines in both settings, but the gains are larger in the setting using neural network function approximation. This is largely due to the relatively poor performance of Deep-FVI and Deep-FQI, which appear to suffer from higher variance and instability during training. We hypothesize that this is due to the difficulty of fitting high-capacity models under long-horizon dynamics, where bootstrapping can amplify noise. In contrast, the structured nature of Deep-FSFVI, specifically the reduced effective discount factor and frozen slow states, likely contributes to improved stability and more reliable learning.
@@
    Again, we see that that the $T=3$ version of Deep-FSFVI performed poorly under both architectures, but $T=6$ and $T=12$ were both adequate.
@@ Conclusion
Appendix A Empirical Versions of VI and FSVI
1
Input: Initial values $U^{0}$, number of iterations $k$, and number of next state samples $M$.
Output: Approximation to the optimal policy $\nu^{k}$.
2 for $k^{\prime}=1,2,\ldots,k$ do
3       for $s$ in the state space $\mathcal{S}$ do
4             $U^{k^{\prime}}(s)=\max_{a}r(s,a)+\gamma\,\frac{1}{M}\sum_{j=1}^{M}\,U^{k^{\prime}-1}(s^{(a,j)})$, where $s^{(a,j)}$ is sampled from the generative model, conditional on state $s$ and action $a$.
5       end for
6
7 end for
8 for $s$ in the state space $\mathcal{S}$ do
9       $\nu^{k}(s)=\operatorname*{arg\,max}_{a}r(s,a)+\gamma\,\frac{1}{M}\sum_{j=1}^{M}\,U^{k}(s^{(a,j)})$, where $s^{(a,j)}$ is sampled from the generative model, conditional on state $s$ and action $a$.
10 end for
11
Algorithm 5 Empirical VI for the Base Model (Base E-VI)
1
Input: Initial values $J_{T}^{*}\equiv 0$ and $V^{0}$, number of iterations $k$, and number of next-state samples $M_{l}$ (lower level), $M_{u}$ (upper level).
2 for each slow state $x\in\mathcal{X}$ do
3       for $t=T-1,T-2,\ldots,1$ do
4             for each fast state $y\in\mathcal{Y}$ do
5                   $J^{*}_{t}(x,y)=\max_{a}r(x,y,a)+\gamma\,\frac{1}{M_{l}}\sum_{j=1}^{M_{l}}J^{*}_{t+1}(x,y^{(a,j)})$, where $y^{(a,j)}$ is sampled from the generative model, conditional on state $(x,y)$ and action $a$.
6
7                   $\tilde{\pi}_{t}^{*}(x,y)=\operatorname*{arg\,max}_{a}\,r(x,y,a)+\gamma\,\frac{1}{M_{l}}\sum_{j=1}^{M_{l}}J^{*}_{t+1}(x,y^{(a,j)})$, with $y^{(a,j)}$ defined the same way as above.
8             end for
9
10       end for
11
12 end for
13for $k^{\prime}=1,2,\ldots,k$ do
14       for $s_{0}=(x_{0},y_{0})$ in the state space $\mathcal{X}\times\mathcal{Y}$ do
15             $V^{k^{\prime}}(x_{0},y_{0},J^{*}_{1},\tilde{\bm{\pi}}^{*})=\max_{a}\,\frac{1}{M_{u}}\sum_{j=1}^{M_{u}}\,\Bigl{[}\tilde{R}^{(a,j)}+\gamma^{T}V^{k^{\prime}-1}(x^{(a,j)}_{T},y^{(a,j)}_{T},J^{*}_{1},\tilde{\bm{\pi}}^{*})\Bigr{]}$, where $x_{T}^{(a,j)}$, $y_{T}^{(a,j)}$, and $\tilde{R}^{(a,j)}$ are sampled using a $T$-step transition through the generative model, conditioned on $x_{0}$, $y_{0}$, $a$, $\tilde{\bm{\pi}}^{*}$ and $J^{*}_{1}$.
16       end for
17
18 end for
19 for $s_{0}=(x_{0},y_{0})$ in the state space $\mathcal{X}\times\mathcal{Y}$ do
20       $\tilde{\mu}^{k}(x_{0},y_{0})=\operatorname*{arg\,max}_{a}\frac{1}{M_{u}}\sum_{j=1}^{M_{u}}\Bigl{[}\tilde{R}^{(a,j)}+\gamma^{T}V^{k}(x^{(a,j)}_{T},y^{(a,j)}_{T},J^{*}_{1},\tilde{\bm{\pi}}^{*})\Bigr{]}$, where $x_{T}^{(a,j)}$, $y_{T}^{(a,j)}$, and $\tilde{R}^{(a,j)}$ are sampled as above, conditioned on $x_{0}$, $y_{0}$, $a$, $\tilde{\bm{\pi}}^{*}$ and $J^{*}_{1}$.
21 end for
22
Algorithm 6 Empirical Frozen-State Value Iteration (E-FSVI)
Appendix B Implementation Details for Experimental Results
Ignoring the Slow State
For Slow-Agn Base VI, our goal is to simulate the situation where a decision maker neglects to consider the slow state during modeling. In other words, we model the problem as if the state consists only of the fast component, intentionally ignoring the slow state altogether. We implement this in our experiments as follows:
  • Full-to-partial state conversion. The environment always operates on the full state space, which includes both fast and slow components. However, the Slow-Agn E-VI algorithm only observes the fast component (i.e., a partial state). Whenever the environment returns a state to the algorithm, it extracts and passes only the fast component.
  • Partial-to-full state conversion. When the algorithm queries the environment (e.g., for rewards or next states) using a partial state and an action, we construct a corresponding full state by randomly sampling a slow state uniformly and appending it to the given fast component. This synthesized full state is then used by the environment to transition forward.
@@ Appendix B Implementation Details for Experimental Results
Appendix C Value Function Approximation with a Linear Architecture
@@
It is well-known that naive specifications of approximate value iteration applied to linear architectures can produce divergent behavior (Bertsekas and Tsitsiklis, 1996). To circumvent this potential issue, our algorithmic approach depends on a set of pre-selected states $\tilde{\mathcal{S}}=\{s_{1},s_{2},\ldots,s_{M}\}$, an idea popularized in Tsitsiklis and Van Roy (1996), who showed that if certain assumptions on these states and the feature vectors are satisfied, then divergence is avoided. A similar algorithm is also described more recently in Zanette et al. (2019).
@@
Appendix D Proofs from Sections 3 and 4
@@
Lemma D.1 is a consequence of Assumption 1. ∎
@@
Lemma D.3 (Discrepancy between $H$ and $\tilde{H}$).
@@
Appendix E Proofs for Section 6 and Section 7
@@
Lemma E.4 (Proposition 6.1 of Bertsekas and Tsitsiklis (1996)).
@@
Noting that $U^{*}$ has Lipschitz constant $L_{U}$ by Assumption 2, we can apply Lemma D.3 to Term A to obtain
Moving on to Term B, since the reward function $r\geq 0$, it follows that $U^{*}(s)\geq 0$ for all $s$. Also, the monotonicity of $\tilde{H}$ implies that $(\tilde{H}^{T-1}U^{*})\geq(\tilde{H}^{T-1}0)$. Therefore, applying Lemma D.2,
$=\gamma^{T-1}\textstyle\max_{y\in\mathcal{Y}_{s_{1}}^{T-1}}U^{*}(x_{1},y),$
$\leq\gamma^{T}L_{U}d_{\mathcal{Y}}(\alpha+2)(T-1),$ (32)
where (31) follows by Assumption 2 and (32) comes from Lemma D.1 (we use that $x_{T}(\omega)$ is $T-1$ transitions from $x_{1}(\omega)$ and both $\tilde{y}_{s_{1}}(\omega)$ and $y_{T}(\omega)$ are both $T-1$ transitions from $y_{1}(\omega)$). Finally, we have
[…]
which completes the proof.
Proof of Lemma 6.1
Let $\epsilon_{r}(\bm{\pi}^{*},J_{1})=\max_{s,a}\,|\mathbb{E}[R(s,a,\bm{\pi}^{*})]-\mathbb{E}[\tilde{R}(s,a,J_{1})]|$, so that we have $|r_{1}(s,a)-r_{2}(s,a)|\leq\epsilon_{r}(\bm{\pi}^{*},J_{1})$. Noting that the first steps of $f_{1}$ and $f_{2}$ are the same (action $a$ in state $s$ with $w_{1}$ revealed), applying parts (ii) and (iii) of Lemma D.1, the maximum discrepancy between $f_{1}$ and $f_{2}$ is:
Applying Lemma E.1, we see that
We also need to account for the fact that $\mu$ is greedy with respect to $V$, an approximation of the optimal value of $\mathcal{M}_{2}$. More precisely, $\mu$ is greedy with respect to $r_{2}(s,a)=\mathbb{E}\bigl{[}\tilde{R}(s_{0},a,J_{1})\bigr{]}$, $f_{2}(s,a,\bm{w})=s_{T}(a,\bm{\pi})$, and value function $V$. We can thus apply Lemma E.2 with $\epsilon_{r}=\epsilon_{r}(\bm{\pi}^{*},J_{1})$, $d=d(\alpha,d_{\mathcal{Y}},T)$, $L_{1}=L_{U}$, $\epsilon_{U}=\bigl{\|}\bar{U}^{*}-V^{*}(J_{1},\bm{\pi})\bigr{\|}_{\infty}$, and $\tilde{\epsilon}_{U}=\bigl{\|}V-V^{*}(J_{1},\tilde{\bm{\pi}})\bigr{\|}_{\infty}$. Collecting terms completes the proof.
@@
Proof of Proposition 7.1
Since $\nu_{k}$ is greedy with respect to $U_{k}$, we can apply Lemmas E.3 and E.4 to obtain
@@ Appendix F Bounds on $L_{U}$ in Terms of $L_{r}$ and $L_{f}$
Appendix G Proofs for Appendix C